can you solve this
for two consecutive odd integers, the sum of the smaller and 4 times the larger is 33. find the integers.

Answer by fractalier(6550)

Let the numbers be
x and x+2
Now translate and get
x + 4(x+2) = 33
x + 4x + 8 = 33
5x + 8 = 33
5x = 25
x = 5
x+2 = 7